OpenStudy (anonymous):

A rechargeable battery in a laptop computer is regularly used over a 24 hour period with the following units consumed. However, the battery only has a charge of 4 units at the beginning of today. The battery will be continually charging at a rate of 1/2 unit per hour. You can assume that the consumption of units is uniform within each 8 hour period. Time period Total units used during the period 00:00 – 08:00 4 08:00 – 16:00 8 16:00 – 24:00 4 When does the battery completely run out of charge? A 04:00 B 08:00 C 12:00 D 16:00 E 24:00

OpenStudy (anonymous):

how much charge does it have at 08:00?

OpenStudy (anonymous):

00:00-08:00=4 units 08:00-16:00=8units 16:00-24:00=4 units

OpenStudy (anonymous):

no, that tells you how much is used in total units between 0:00 and 8:00, not how full the battery is

OpenStudy (anonymous):

to find the charge at 01:00, you need to take away any charge used, and add any charge recharged. For example the battery consumes 4/8 = 1/2 during each of the first 8 hours, and also replenishes 1/2 during each of the first 8 hours. But during 08:00 to 16:00 it uses up more charge.

OpenStudy (anonymous):

ok thank you